Five electrons have been removed from each of two atoms to form ions. These ions are placed in free space with a distance 20 angstrom between them. Find the electrostatic force between them.

Grade:12

2 Answers

chaman
10 years ago
F=K(Q/(2×10-9)-4×10-10) WHERE Q=TOTAL CHARGE OF ATOM BEFORE REMOVING OF ELECTRON
Swathi
7 years ago
F=K × q1 q2 /r^2
q1=q2=q,q=ne.  .....Q=5×1.6×10^-19=8×10^-19
F=9×10^9×8×10^-19×8×10^-19/20×20×10^-10×10^-10=1.44nN